Find the force between two point charges 5uc and 10uc separated by a distance of 0.2m in air.


1
Expert's answer
2021-03-18T14:03:59-0400
"F=\dfrac{kq_1q_2}{r^2},""F=\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot5\cdot10^{-6}\ C\cdot10\cdot10^{-6}\ C}{(0.2\ m)^2}=11.25\ N."